40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 20 13:59:30 2018, Koji, Update, General, Modulation depth measurement for an aLIGO EOM 
    Reply  Thu Feb 22 16:11:14 2018, Koji, Update, General, aLIGO EOM crystal replacement IMG_3421.JPGIMG_3426.JPGIMG_3427.JPG
       Reply  Thu Feb 22 17:19:47 2018, Koji, Update, General, Modulation depth measurement for an aLIGO EOM IMG_3436.JPGmodulation_depth.pdfmodulation_linearity.pdfmodulation_24MHz.pdf
          Reply  Mon Apr 2 15:14:21 2018, Koji, Update, General, Modulation depth measurement for an aLIGO EOM modulation_depth.pdfimpedance_eom.pdf
Message ID: 13642     Entry time: Tue Feb 20 13:59:30 2018     Reply to this: 13650
Author: Koji 
Type: Update 
Category: General 
Subject: Modulation depth measurement for an aLIGO EOM 

Last night I worked at the PSL table for the modulation depth measurement for an aLIGO EOM. Let me know if the IFO behavior is unusual.

What I did was:

  • Cranked up the HEPA speed to 100
  • Placed an aLIGO EOM in the AUX beat path (south side of the PSL laser). (It is still on the PSL table as of Feb 20, 2018)
  • Closed the PSL shutter
  • Turned off the main Marconi forr 11MHz. The freq and output power of this marconi have not been touched.
  • Turned off the freq generation unit
     
  • Worked on my measurement with the spectrum and network analyzers + aux marconi.
     
  • Turned down the HEPA speed to 30
  • Turned on the freq generation unit
  • Turned on the main Marconi
  • Opened the PSL shutter => IMC locked
ELOG V3.1.3-